OpenStudy (anonymous):

16 is 40% of what number

OpenStudy (deadpoolroxz):

Well, 16/x=40/100 100(16)=x(40) 1600=x(40) x=40

OpenStudy (deadpoolroxz):

dat help? if so medel be mine

OpenStudy (anonymous):

whats the percent of change between 14 inches to 35 inches

OpenStudy (deadpoolroxz):

before i give anser google it 4real

OpenStudy (anonymous):

already did

OpenStudy (deadpoolroxz):

also put in open study or brainly on it

OpenStudy (deadpoolroxz):

168

OpenStudy (deadpoolroxz):

got it?

OpenStudy (deadpoolroxz):

@sleepyhead314

OpenStudy (deadpoolroxz):

@logan13

OpenStudy (deadpoolroxz):

@bohotness

OpenStudy (deadpoolroxz):

amiright?

OpenStudy (deadpoolroxz):

helo? u dere @edwards555

OpenStudy (anonymous):

ya did it